10 kids are arranging themselves in 3 different seats(in order). How many ways can they arrange themselves?
Answer by stanbon(75887)

10 kids are arranging themselves in 3 different seats (in order). How many ways can they arrange themselves?
--------
Ans: 10P3 = 10!/(10-3)! = 10!/7! = 10*9*8 = 720
=============
Cheers,
Stan H.
=============